How many liters of pure water must be added to 64 liters of 12% salt solution to reduce it to a 10% salt solution?
Answer by ewatrrr(24785)


Hi
.12*64L = 7.68 salt
.10*76.8 = 7.68 salt
76.8L - 64L = 12.8L of water need to be added